A moving coil instrument has a full scale deflection of 20mA and a resistance of 25ohms , determine the value of the resistor to be used as a,,, shunt to enable the instrument to be used as a 0-20AAmmeter,and multiplier to enable the instrument to be used as a 0-20v voltmeter

1
Expert's answer
2022-02-10T02:15:03-0500

(i) The resistance is

r=rqIq/I-Ia = 25*20*10-3/20-20*10-3 = 0.025Ω.

(ii) The resistance is

r=rv(V/Vv-1)

r=25(20/25*0.02-1)

r=25(40-1)

r=975Ω.